OpenStudy (anonymous):

Simplify the expression. {[1/(1+x+h)]+[1/(x+1)]} / h

OpenStudy (unklerhaukus):

\[ \frac{{\left[\frac{1}{1+x+h}\right]+\left[\frac1{x+1}\right]}}h\]

OpenStudy (anonymous):

how do you simplify that?

OpenStudy (anonymous):

take LCM :) then solve....

OpenStudy (anonymous):